# Scalar-Induced Gravitational Waves from self-resonant preheating in $\bm{\alpha}$-attractor models

### v1

@@ 3.2 Numerical strategy
In Inomata:2019zqy ; Inomata:2019ivs , it is pointed out that the spectrum of SIGWs corresponding to the modes $k<k_{\rm end}$ could be affected by the later transition from the preheating (inflaton matter domination) to the radiation era. In the case of a gradual transition from preheating to radiation, Ref. Inomata:2019zqy indicates that the SIGWs enhancement gets suppressed because of the subsequent decay of the gravitational potential $\Phi_{k}$. However, in the case of a sudden transition Inomata:2019ivs , the same conclusion does not apply; in fact, there is an additional enhancement in SIGW. By comparison to Inomata:2019zqy ; Inomata:2019ivs , in this paper, we also consider in addition the self-resonance effects in the $k>k_{\rm end}$ region of the spectrum (i.e., the modes that never exit the horizon during inflation but fall into the instability band). Notably the enhanced contribution to SIGW $\Omega_{\text{GW}}^{\text{TOT}}$ is dominated by the modes $k\gg k_{\rm end}$ as we can notice from Fig. 4 and Fig. 5. A recent study of this regime Chakraborty:2025zgx shows the gravitational particle production contribution from the small wavelength modes $k\gg k_{\rm end}$ is negligible because the vacuum associated with these modes during inflation is closer to the Minkowski vacuum rather than the de Sitter vacuum. This indicates that the decay of gravitational potential $\Phi_{k}$ for these modes is unlikely to happen. However, a precise analytical and numerical study of this scenario is left for future work.
4 Conclusions
@@
In this work, we explored the production of Scalar-Induced Gravitational Waves (SIGWs) during the preheating phase in the context of $\alpha$-attractor inflationary models. These models, characterized by the parameter $\alpha$, deviate from a quadratic potential as $\alpha$ decreases (see Fig. 3), which triggers a self-resonance effect and rapidly and substantially amplifies the scalar perturbations during preheating del-Corral:2024vcm ; Hertzberg:2014iza ; Hertzberg:2014jza . We particularly considered all the perturbations that fall into the preheating instability band, which include both kinds of modes that exit and never exit the horizon during inflation, i.e., $k<k_{\rm end}$ and $k>k_{\rm end}$. In particular, we found that the modes $k>k_{\rm end}$ whose wavelength is larger than the Jeans length are significantly amplified, and these modes would not contribute to particle production in the post-preheating phase. This amplification, in turn, directly impacts the production of gravitational waves, since scalar perturbations couple to the tensor perturbations at second order through the source term (See  (12) and (13)). We computed the total fractional energy density of GWs (SIGWs + BGWs) for different values of $\alpha$ and our findings show that, for sufficiently small values of $\alpha$, the integral of the total fractional energy density of GWs reaches the BBN bound (32), see Figs. 5 and 6. In particular, we observe that the growth in $\Omega_{GW}^{\text{TOT}}$ is extremely large (by several orders of magnitude) for small differences in $\alpha$. See for instance Fig. 5 for $\alpha=10^{-2}-10^{-3}$. This is due to the fact that the fractional energy density of SIGWs is, naively speaking, proportional to $\int\mathcal{P}_{\Phi}^{2}(k)\sim\int\mathcal{P}_{\mathcal{R}}^{2}(k)$ (See  (21)), meaning small variations in $\mathcal{R}_{\bm{k}}$ induce large changes in $\Omega_{GW}^{\text{SI}}$ so that the secondary effect (SIGWs) dominates over the linear effect (BGWs). In contrast, this amplification mechanism is suppressed for higher values of $\alpha$, including the Starobinsky case, where the effect becomes negligible.
@@
We demonstrated, for the first time, that the lower limit on $\alpha$ imposed by the BBN constraint differs between T-models and E-models when considering the production of SIGWs during preheating instabilities. This approach relies on the assumption that the decay of the inflaton field is not instantaneous, with its coupling to the fields responsible for reheating the universe being sufficiently small to ensure that the preheating phase lasts at least 5 e-folds before the inflaton decays into ultra-relativistic particles Kofman:1994rk ; Kofman:1997yn .
@@
The distinction in the lower bound on $\alpha$ arises from the additional amplification present in the E-model compared to the T-model, which can be attributed to the asymmetry of the potential del-Corral:2024vcm . These new bounds, although derived under different assumptions, are approximately one order of magnitude above those reported in Iacconi:2023mnw , where $\log_{10}(\alpha)=-4.2^{+5.4}_{-8.6}$ at $95\%$ C.L. Furthermore, they establish a new lower bound on the tensor-to-scalar ratio, as given in (33).
@@
In conclusion, our analysis suggests that the amplification of scalar perturbations during self-resonant preheating offers an effective mechanism for generating GWs, particularly in the VHF band. This remarks the need for increasing the sensitivity of gravitational wave detectors in the VHF range that could potentially test these theoretical predictions Aggarwal:2020olq ; Kuroda:2015owv ; Page:2020zbr ; Dandoy:2024oqg and offer new insights into the physics of the early universe. Furthermore, we derive for the first time a concrete lower bound on the tensor-to-scalar ratio for the single-field Starobinsky-like inflationary scenarios, which would be an important input for the future probes of B-mode polarization Belkner:2023duz ; CMB-S4:2016ple .
@@ Acknowledgements.
D.C. is grateful for the support of grant UI/BD/151491/2021 from the Portuguese Agency Fundação para a Ciência e a Tecnologia. The work of P.G. was partially supported by NSF grant PHY-2412829. P.G. thanks Prof. Teruaki Suyama for his kind hospitality at the Institute of Science Tokyo. This research was funded by Fundação para a Ciência e a Tecnologia grant number UIDB/MAT/00212/2020 and COST action 23130. K.S.K. acknowledges the support from the Royal Society Newton International Fellowship. We would like to thank David Wands for very useful discussions.
@@ Appendix A Source term for SIGWs
In this appendix, we closely follow the analysis presented in Domenech:2021 and Baumann:2007 . By considering both metric and field fluctuations, we can write the following line element for the perturbed FLRW metric
$\mathrm{d}s^{2}=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}=-(1+2\Psi)\mathrm{d}t^{2}+a^{2}[(1-2\Phi)\delta_{ij}+h_{ij}]\mathrm{d}x^{i}\mathrm{d}x^{j},$ (34)
where $h_{ij}$ is a gauge-invariant, symmetric, transverse, and traceless perturbation tensor. We are working in the Newtonian gauge. The most general form for $h_{ij}$ is given by
@@
where $\tau_{ij}^{(\lambda)}$ are polarization tensors, which are symmetric, transverse, and traceless, and the index $\lambda=+,\times$ denotes the two polarization states. Using the metric defined in (34) we can derive the following field equation for $h_{ij}$ from the third order perturbative expansion of the action (4) Guzzetti:2016 ; Bauman:2009
@@
Here $\mathcal{S}_{lm}$ is the source term, which contains the information about scalar perturbations and is given by Domenech:2021
@@
The term $\mathcal{T}_{ij}^{lm}$ is the projection tensor, responsible for extracting the transverse and traceless components of any tensor and the terms involving second-order scalar and tensor perturbations Ananda:2006 . It can be expressed in terms of the polarization tensors $\tau_{ij}^{(\lambda)}$ as Baumann:2007
$\mathcal{T}_{ij}^{lm}\mathcal{S}_{lm}=\int\frac{\mathrm{d}^{3}\mathbf{k}}{(2\pi)^{3/2}}e^{i\mathbf{k}\cdot\mathbf{x}}\left(\sum_{\lambda=+,\times}\tau_{ij}^{(\lambda)}(\mathbf{k})\tau^{(\lambda)lm}(\mathbf{k})\right)\mathcal{S}_{lm}(\mathbf{k}),$ (38)
@@
By applying eqns. (38) and (39), we can transform the equation of motion (36) into Fourier space, resulting in the following expression Baumann:2007
@@
Next, we compute the Fourier transform of $\mathcal{S}_{lm}(\bm{x})$, for which we use the following property of the Fourier transforms, called the Convolution Theorem. Consider two functions, $f$ and $g$. The Fourier transform of their product is given by Arfken:379118
@@
$(f\ast g)(\bm{k})=\int\mathrm{d}^{3}\tilde{\bm{k}}\,f(\tilde{\bm{k}})\,g(\bm{k}-\tilde{\bm{k}}).$ (42)
Using this, the source term in (40) is now given by Baumann:2007 ; Ananda:2006 ; Assadullahi:2009
$\begin{split}\mathcal{S}(\bm{k},t)&=-\tau^{lm}\mathcal{S}_{lm}(\bm{k},t)\\ &=\int\frac{\mathrm{d}^{3}\tilde{k}}{(2\pi)^{3/2}}\tau^{lm}\tilde{k}_{l}\tilde{k}_{m}\left(4\Phi_{\tilde{\bm{k}}}\Phi_{\bm{k}-\tilde{\bm{k}}}+2\delta\phi_{\tilde{\bm{k}}}\delta\phi_{\bm{k}-\tilde{\bm{k}}}\right),\end{split}$ (43)
where we are ignoring the anisotropic stress, which translates into $\Phi+\Psi=0$. The meaning of (43) is that, since the source term consists of products of perturbations that interact in real space and thus their Fourier coefficients mix, the resulting mode of the product is influenced by all possible pairs of modes of the original perturbations, which translates into a convolution in $\bm{k}$. Using the properties of the projection tensors, we have that Baumann:2007
@@
$\mathcal{S}(\bm{k},t)=\int\frac{\mathrm{d}^{3}\tilde{k}}{(2\pi)^{3/2}}\tilde{k}^{2}(1-\mu^{2})\left(4\Phi_{\tilde{\bm{k}}}\Phi_{\bm{k}-\tilde{\bm{k}}}+2\delta\phi_{\tilde{\bm{k}}}\delta\phi_{\bm{k}-\tilde{\bm{k}}}\right).$ (45)
@@ Appendix B Power spectrum of SIGWs
As shown in del-Corral:2024vcm , the self-resonance is brief for small values of $\alpha$, followed by the usual, pressureless, matter-dominated preheating, where the scale factor evolves as $a\sim t^{2/3}$ and $\Phi_{\bm{k}}\sim\delta\phi_{\bm{k}}\sim C(k)$ for the modes of interest. We remark that this does not imply any fluid description since the matter content of the universe is still in the form of an oscillating scalar field. Under these assumptions, $\mathcal{S}(\bm{k},t)=\mathcal{S}_{\bm{k}}$, and we can attempt to solve (12) semi-analytically. A particular solution is given by Assadullahi:2009 :
@@
since the typical assumption is that SIGWs are generated instantaneously when the relevant mode enters the horizon Baumann:2007 . Using (46), the two-point correlation function for $h_{\bm{k}}$ can be written as
$\begin{split}\langle h_{\bm{k}}^{(\lambda)}(t)h_{\bm{k}^{\prime}}^{(\lambda)}(t)\rangle&=\frac{g^{2}(k,t)}{k^{4}}\langle\mathcal{S}_{\bm{k}}\mathcal{S}_{\bm{k}^{\prime}}\rangle\\ &\simeq\frac{16g^{2}(k,t)}{k^{4}}\int\frac{\mathrm{d}^{3}\tilde{\bm{k}}}{(2\pi)^{3/2}}\int\frac{\mathrm{d}^{3}\tilde{\bm{k}}^{\prime}}{(2\pi)^{3/2}}\tilde{k}^{2}\tilde{k}^{\prime 2}(1-\mu^{2})^{2}\langle\Phi_{\tilde{\bm{k}}}\Phi_{\bm{k}-\tilde{\bm{k}}}\Phi_{\tilde{\bm{k}^{\prime}}}\Phi_{\bm{k}-\tilde{\bm{k}^{\prime}}}\rangle\\ &\simeq\frac{8\pi g^{2}(k,t)}{k^{4}}\delta^{(3)}(\bm{k}+\bm{k}^{\prime})\int\mathrm{d}^{3}\tilde{\bm{k}}\,\frac{\tilde{k}\,(1-\mu^{2})^{2}}{|\bm{k}-\tilde{\bm{k}}|^{3}}\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\phi}(|\bm{k}-\tilde{\bm{k}}|),\end{split}$ (48)
where in the second step we are neglecting $\delta\phi_{\bm{k}}$ since it is subdominant with respect to $\Phi_{\bm{k}}$, and in the last equality we used Wick’s theorem to express the four-point correlation function of a Gaussian distribution of scalar perturbations as a function of two-point correlation functions Assadullahi:2009 ; Baumann:2007 ; Domenech:2021 . Using the definition of the power spectrum of SIGWs, eqn. (15), we have
$\mathcal{P}^{\text{SI}}_{h}(k,t)=\frac{8g^{2}(k,t)}{\pi k}\int\mathrm{d}^{3}\tilde{\bm{k}}\,\frac{\tilde{k}\,(1-\mu^{2})^{2}}{|\bm{k}-\tilde{\bm{k}}|^{3}}\,\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\Phi}(|\bm{k}-\tilde{\bm{k}}|).$ (49)
@@
$\begin{split}\int_{V}\mathrm{d}^{3}\tilde{\bm{k}}&=\int_{0}^{\infty}\tilde{k}^{2}\mathrm{d}\tilde{k}\int_{0}^{\pi}\sin{\theta}\mathrm{d}\theta\int_{0}^{2\pi}\mathrm{d}\varphi=\int_{0}^{\infty}\tilde{k}^{2}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\int_{0}^{2\pi}\mathrm{d}\varphi,\end{split}$ (50)
@@
$\begin{split}\mathcal{P}_{h}^{\text{SI}}(k,t)&=\frac{16g^{2}(k,t)}{k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\,\frac{\tilde{k}^{3}\,(1-\mu^{2})^{2}}{|\bm{k}-\tilde{\bm{k}}|^{3}}\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\Phi}(|\bm{k}-\tilde{\bm{k}}|).\end{split}$ (51)
@@
$\mathcal{P}^{\text{SI}}_{h}(k,t)\sim\frac{256g^{2}(k,t)}{15k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\,\mathcal{P}^{2}_{\Phi}(\tilde{k})=\frac{256g^{2}(k,t)\,\mathcal{I}_{\tilde{k}}}{15k},$ (52)
where $\mathcal{I}_{\tilde{k}}$ represents the integral of the square of the power spectrum over the whole range of $\tilde{k}$. This essentially depends on the amplitude of the peak, so the higher the peak, the higher the integral, which ultimately translates into an amplification for the low-$k$ range, although the peak is in the high-$k$ range. This clearly reflects the convolutive nature of the source term, as stated in Sec. 3.1. Regarding the growth function $g(k,t)$, let us write $x=\frac{2k}{aH}$. Then, the expansion of $g(x)$ for small $x$ is $g\sim x^{2}/10$ and therefore, eqn. (52), evaluated at the end of preheating, is given by
@@
$\mathcal{P}_{h}^{\text{SI}}(k\gg\tilde{k},t_{rh})\sim\frac{16\,\mathcal{P}_{\Phi}(k)}{k^{4}}\int_{0}^{\infty}\tilde{k}^{3}\mathcal{P}_{\Phi}(\tilde{k})=\frac{16\,\mathcal{P}_{\Phi}(k)\mathcal{I}_{k}}{k^{4}},$ (56)
where $\mathcal{I}_{k}$ is defined in a similar way to $\mathcal{I}_{\tilde{k}}$ and the growth function is set to unity. This translates into the following expression for the fractional energy density
$\Omega_{\text{GW}}^{\text{SI}}(k\gg\tilde{k},t_{rh})\sim\frac{4\Omega_{\gamma}^{0}\,\mathcal{P}_{\Phi}(k)\mathcal{I}_{k}}{3k_{U}^{2}k^{2}}$ (57)
@@
where $k_{p}\sim k_{U}/\sqrt{2}$ Hertzberg:2014iza ; Hertzberg:2014jza . Eqn. (58) is shown as the vertical blue dashed line of Fig. 7. On the contrary, for $\alpha=1$, since there is no amplification peak, the maximum is at the point where the growth function peaks. Again, considering $x=\frac{2k}{aH}$, these points are given by the roots of the equation

### v2

@@ III.2 Numerical strategy
In [81, 82], it is pointed out that the spectrum of SIGWs corresponding to the modes $k<k_{\rm end}$ could be affected by the later transition from the preheating (inflaton matter domination) to the radiation era. In the case of a gradual transition from preheating to radiation, Ref. [81] indicates that the SIGWs enhancement gets suppressed because of the subsequent decay of the gravitational potential $\Phi_{k}$. However, in the case of a sudden transition [82], the same conclusion does not apply; in fact, there is an additional enhancement in SIGW. By comparison to [81, 82], in this paper, we also consider in addition the self-resonance effects in the $k>k_{\rm end}$ region of the spectrum (i.e., the modes that never exit the horizon during inflation but fall into the instability band). Notably the enhanced contribution to SIGW $\Omega_{\text{GW}}^{\text{TOT}}$ is dominated by the modes $k\gg k_{\rm end}$ as we can notice from Fig. 4 and Fig. 5. A recent study of this regime [83] shows the gravitational particle production contribution from the small wavelength modes $k\gg k_{\rm end}$ is negligible because the vacuum associated with these modes during inflation is closer to the Minkowski vacuum rather than the de Sitter vacuum. This indicates that the decay of gravitational potential $\Phi_{k}$ for these modes is unlikely to happen. However, a precise analytical and numerical study of this scenario is left for future work.
IV Conclusions
@@
In this work, we explored the production of Scalar-Induced Gravitational Waves (SIGWs) during the preheating phase in the context of $\alpha$-attractor inflationary models. These models, characterized by the parameter $\alpha$, deviate from a quadratic potential as $\alpha$ decreases (see Fig. 3), which triggers a self-resonance effect and rapidly and substantially amplifies the scalar perturbations during preheating [23, 24, 25]. We particularly considered all the perturbations that fall into the preheating instability band, which include both kinds of modes that exit and never exit the horizon during inflation, i.e., $k<k_{\rm end}$ and $k>k_{\rm end}$. In particular, we found that the modes $k>k_{\rm end}$ whose wavelength is larger than the Jeans length are significantly amplified, and these modes would not contribute to particle production in the post-preheating phase. This amplification, in turn, directly impacts the production of gravitational waves, since scalar perturbations couple to the tensor perturbations at second order through the source term (See  (14) and (15)). We computed the total fractional energy density of GWs (SIGWs + BGWs) for different values of $\alpha$ and durations of preheating ($\Delta N$), and our findings show that, for sufficiently small values of $\alpha$, the integral of the total fractional energy density of GWs reaches the BBN bound (34), see Figs. 5 and 6. In particular, we observe that the growth in $\Omega_{GW}^{\text{TOT}}$ is extremely large (by several orders of magnitude) for small differences in $\alpha$. See for instance Fig. 5 for $\alpha=10^{-2}-10^{-3}$. This is due to the fact that the fractional energy density of SIGWs is, naively speaking, proportional to $\int\mathcal{P}_{\Phi}^{2}(k)\sim\int\mathcal{P}_{\delta\phi}^{2}(k)$ (See  (24)), meaning small variations in the perturbations induce large changes in $\Omega_{GW}^{\text{SI}}$ so that the secondary effect (SIGWs) dominates over the linear effect (BGWs). In contrast, this amplification mechanism is suppressed for higher values of $\alpha$, including the Starobinsky case, where the effect becomes negligible.
@@
We demonstrated, for the first time, that the lower limit on $\alpha$ imposed by the BBN constraint differs between T-models and E-models when considering the production of SIGWs during preheating instabilities. This approach relies on the assumption that the decay of the inflaton field is not instantaneous, with its coupling to the fields responsible for reheating the universe being sufficiently small to ensure that the preheating phase lasts at least 1 or 2 e-folds before the inflaton decays into ultra-relativistic particles [18, 19]. We also stress the fact that the transition into the standard radiation-dominated universe after the inflaton’s decay into particles is not studied in this paper. However, as pointed out by the end of Section III.2, it is unlikely that this could affect our results. In addition, we note that our derived bounds on tensor-to-scalar ratio are obtained at the level of linearised tensor fluctuations that include the effect of scalar-scalar-tensor interactions with metric and scalar field matter fluctuations. However, our results are subject to verification with the complete non-linear analysis, including all metric fluctuations at higher orders, which would require techniques beyond the existing frameworks of lattice simulations [84, 85]. We defer this investigation for future studies.
@@
The distinction in the lower bound on $\alpha$ (for the same $\Delta N$) arises from the additional amplification present in the E-model compared to the T-model, which can be attributed to the asymmetry of the potential [23]. These new bounds, although derived under different assumptions, should be compared with those reported in [10] based on the duration of reheating along with the current bounds on the scalar spectral index, where $\log_{10}(\alpha)=-4.2^{+5.4}_{-8.6}$ at $95\%$ C.L, in [86] based on the overproduction of a light moduli field, where $\alpha\lesssim 10^{-8}$, or in [87] based on geometrical destabilization of a spectator field, where $\alpha\lesssim 10^{-3}$. Furthermore, they establish a new lower bound on the tensor-to-scalar ratio, as given in Table 1.
@@
In conclusion, our analysis suggests that the amplification of scalar perturbations during self-resonant preheating offers an effective mechanism for generating GWs, particularly in the VHF band. This remarks the need for increasing the sensitivity of gravitational wave detectors in the VHF range that could potentially test these theoretical predictions [46, 45, 47, 48] and offer new insights into the physics of the early universe. Furthermore, we derive for the first time a concrete lower bound on the tensor-to-scalar ratio for the single-field Starobinsky-like inflationary scenarios, which would be an important input for the future probes of B-mode polarization [88, 89].
@@ Acknowledgements.
We thank the anonymous referee for careful reading and important suggestions to improve the quality of the manuscript. D.C. is grateful for the support of grant UI/BD/151491/2021 from the Portuguese Agency Fundação para a Ciência e a Tecnologia. The work of P.G. was partially supported by NSF grant PHY-2412829. P.G. thanks Prof. Teruaki Suyama for his kind hospitality at the Institute of Science Tokyo. This research was funded by Fundação para a Ciência e a Tecnologia grant number UIDB/MAT/00212/2020 and COST action 23130. K.S.K. acknowledges the support from the Royal Society Newton International Fellowship. We would like to thank David Wands for the very useful discussions.
@@ Appendix A Source term for SIGWs
In this appendix, we closely follow the analysis presented in [41] and [38]. By considering both metric and field fluctuations, we can write the following line element for the perturbed FLRW metric
$\begin{split}\mathrm{d}s^{2}&=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}\\ &=-(1+2\Phi)\mathrm{d}t^{2}+a^{2}[(1-2\Psi)\delta_{ij}+h_{ij}]\mathrm{d}x^{i}\mathrm{d}x^{j},\end{split}$ (35)
where $h_{ij}$ is a gauge-invariant, symmetric, transverse, and traceless perturbation tensor. We are working in the Newtonian gauge and we impose $\Phi=\Psi$ constraint since there is no anisotropic stress involved for the scalar field matter. Thus, the perturbed Einstein equations are given by [90, 91]
$3H\left(\dot{\Phi}+H\Phi\right)+\frac{\nabla^{2}\Phi}{a^{2}}=\frac{1}{M_{\mathrm{Pl}}^{2}}\,\delta\rho\,.$ (36)
where
$\delta T^{0}{}_{0}=-\delta\rho=-\left(\dot{\bar{\phi}}\,\delta\dot{\phi}-\Phi\,\dot{\bar{\phi}}^{2}+V^{\prime}(\bar{\phi})\delta\phi\right)$} (37)
and
$\dot{\Phi}+H\Phi=\frac{1}{2M_{\mathrm{Pl}}^{2}}\,\dot{\bar{\phi}}\,\delta\phi$ (38)
The perturbed scalar field equations of motion are
$\delta\phi^{\prime\prime}+2\mathcal{H}\delta\phi^{\prime}-\nabla^{2}\delta\phi+a^{2}V_{,\phi\phi}\delta\phi-4\bar{\phi}^{\prime}\Phi^{\prime}+2a^{2}V_{,\phi}\Phi=0$ (39)
We can notice clearly from (38) that the metric and scalar fluctuations are always related to each other. In Fig. 8, we can see in detail that, during inflation, the amplitude of metric fluctuations would be a lot smaller than the scalar-field fluctuations, while during preheating instability, both the metric and scalar field fluctuations would get amplified in the same manner due to the resonant instability.
The most general form for $h_{ij}$ is given by
@@
where $\tau_{ij}^{(\lambda)}$ are polarization tensors, which are symmetric, transverse, and traceless, and the index $\lambda=+,\times$ denotes the two polarization states. Using the metric defined in (35) we can derive the following field equation for $h_{ij}$ from the third order perturbative expansion of the action (3) [39, 91]
@@
Here $\mathcal{S}_{lm}$ is the source term, which contains the information about scalar perturbations and is given by [41]
@@
The term $\mathcal{T}_{ij}^{lm}$ is the projection tensor, responsible for extracting the transverse and traceless components of any tensor and the terms involving second-order scalar and tensor perturbations [36]. It can be expressed in terms of the polarization tensors $\tau_{ij}^{(\lambda)}$ as [38]
$\begin{split}\mathcal{T}_{ij}^{lm}\mathcal{S}_{lm}&=\int\frac{\mathrm{d}^{3}\mathbf{k}}{(2\pi)^{3/2}}e^{i\mathbf{k}\cdot\mathbf{x}}\\ &\times\left(\sum_{\lambda=+,\times}\tau_{ij}^{(\lambda)}(\mathbf{k})\tau^{(\lambda)lm}(\mathbf{k})\right)\mathcal{S}_{lm}(\mathbf{k}),\end{split}$ (43)
@@
By applying eqns. (43) and (44), we can transform the equation of motion (41) into Fourier space, resulting in the following expression [38]
@@
Next, we compute the Fourier transform of $\mathcal{S}_{lm}(\bm{x})$, for which we use the following property of the Fourier transforms, called the Convolution Theorem. Consider two functions, $f$ and $g$. The Fourier transform of their product is given by [92]
@@
[…]
Using this, the source term in (45) is now given by [38, 36, 37]
[…]
The meaning of (48) is that, since the source term consists of products of perturbations that interact in real space and thus their Fourier coefficients mix, the resulting mode of the product is influenced by all possible pairs of modes of the original perturbations, which translates into a convolution in $\bm{k}$. Using the properties of the projection tensors, we have that [38]
@@
[…]
@@ Appendix B Power spectrum of SIGWs
As shown in [23], the self-resonance is brief for small values of $\alpha$, followed by the usual, pressureless, matter-dominated preheating, where the scale factor evolves as $a\sim t^{2/3}$. We remark that this does not imply any fluid description since the matter content of the universe is still in the form of an oscillating scalar field. Also after the short self-resonance phase, the perturbations $\Phi_{\bm{k}}$ and $\delta\phi_{\bm{k}}$ evolve approximately constant in amplitude for the modes of interest. Under these assumptions, the source term remains also constant in amplitude and therefore $\mathcal{S}(\bm{k},t)=\mathcal{S}_{\bm{k}}$, and we can attempt to solve (14) semi-analytically. A particular solution is given by [37]:
@@
since the typical assumption is that SIGWs are generated instantaneously when the relevant mode enters the horizon [38]. Using (51), the two-point correlation function for $h_{\bm{k}}$ can be written as
[…]
where in the last equality we used Wick’s theorem to express the four-point correlation function of a Gaussian distribution of scalar perturbations as a function of two-point correlation functions [37, 38, 41]. Using the definition of the power spectrum of SIGWs, eqn. (18), we have
[…]
@@
$\int_{V}\mathrm{d}^{3}\tilde{\bm{k}}=\int_{0}^{\infty}\tilde{k}^{2}\mathrm{d}\tilde{k}\int_{0}^{\pi}\sin{\theta}\mathrm{d}\theta\int_{0}^{2\pi}\mathrm{d}\varphi=\int_{0}^{\infty}\tilde{k}^{2}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\int_{0}^{2\pi}\mathrm{d}\varphi,$ (55)
@@
[…]
@@
$\mathcal{P}^{\text{SI}}_{h}(k,t)\sim\frac{256g^{2}(k,t)}{15k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\bigg[\mathcal{P}^{2}_{\Phi}(\tilde{k})+\frac{1}{2}\mathcal{P}^{2}_{\delta\phi}(\tilde{k})\bigg]=\frac{256g^{2}(k,t)\,\mathcal{I}_{\tilde{k}}}{15k},$ (57)
where $\mathcal{I}_{\tilde{k}}$ represents the integral of the square of the power spectra over the whole range of $\tilde{k}$. This essentially depends on the amplitude of the peak, so the higher the peak, the higher the integral, which ultimately translates into an amplification for the low-$k$ range, although the peak is in the high-$k$ range. This clearly reflects the convolutive nature of the source term, as stated in Sec. III.1. Regarding the growth function $g(k,t)$, let us write $x=\frac{2k}{aH}$. Then, the expansion of $g(x)$ for small $x$ is $g\sim x^{2}/10$ and therefore, eqn. (57), evaluated at the end of preheating, is given by
@@
$\begin{split}\mathcal{P}_{h}^{\text{SI}}(k\gg\tilde{k},t_{rh})&\sim\frac{16\,\mathcal{P}_{\Phi}(k)}{k^{4}}\int_{0}^{\infty}\tilde{k}^{3}\mathcal{P}_{\Phi}(\tilde{k})\\ &+\frac{8\,\mathcal{P}_{\delta\phi}(k)}{k^{4}}\int_{0}^{\infty}\tilde{k}^{3}\mathcal{P}_{\delta\phi}(\tilde{k})\\ &=\frac{16\,\mathcal{P}_{\Phi}(k)\mathcal{I}_{k}^{\Phi}}{k^{4}}+\frac{8\,\mathcal{P}_{\delta\phi}(k)\mathcal{I}_{k}^{\delta\phi}}{k^{4}},\end{split}$ (61)
where $\mathcal{I}_{k}^{\Phi}$ and $\mathcal{I}_{k}^{\delta\phi}$ are defined in a similar way to $\mathcal{I}_{\tilde{k}}$ and the growth function is set to unity. This translates into the following expression for the fractional energy density
$\Omega_{\text{GW}}^{\text{SI}}(k\gg\tilde{k},t_{rh})\sim\frac{4\Omega_{\gamma}^{0}\,\mathcal{P}_{\Phi}(k)\mathcal{I}_{k}^{\Phi}}{3k_{U}^{2}k^{2}}+\frac{2\Omega_{\gamma}^{0}\,\mathcal{P}_{\delta\phi}(k)\mathcal{I}_{k}^{\delta\phi}}{3k_{U}^{2}k^{2}}$ (62)
@@
where $k_{p}\sim k_{U}/\sqrt{2}$ [24, 25]. Eqn. (63) is shown as the vertical blue dashed line of Fig. 7. On the contrary, for $\alpha=1$, since there is no amplification peak, the maximum is at the point where the growth function peaks. Again, considering $x=\frac{2k}{aH}$, these points are given by the roots of the equation